OpenStudy (anonymous):

How do I write An equation in point slope form for this: A train travels at a rate of 70 mph. Two hours after leaving the station it is 210 miles from it's destination.

OpenStudy (anonymous):

???

OpenStudy (chris):

the question gives you the slope - 70 , and one example to find the constant (b)

OpenStudy (chris):

y = 70x +b

OpenStudy (chris):

210 = 70*2 + b

OpenStudy (chris):

b = 210 - 140 = 70

OpenStudy (chris):

y = 70x + 70

OpenStudy (anonymous):

Thanks Chris

OpenStudy (chris):

np man

OpenStudy (anonymous):

YOU DA MAN!